Jenna and Michael are throwing a 0.420 kg football back and forth. Jenna also has a 1.20 kg ball of clay. Michael throws the football horizontally toward Jenna at 3.20 m/s, at the same time Jenna throws the ball of clay toward the football. If the football stops moving forward after the clay impacts the football, how fast did Jenna throw the clay?
1
Expert's answer
2018-10-25T15:43:09-0400

The momentum of a body is given by

p=mv,

where m – is the mass of the body and v – is its velocity.

Since it is given that the mass of the football is m_f=0.42 kg and its speed is v_f=3.2 m/s, for its momentum we obtain

p_f=m_f v_f=0.42 kg⋅3.2 m/s=1.344 (kg⋅m)/s

Let the speed of the clay ball be v_c, then its momentum is

p_c=m_c v_c,

where m_c=1.2 kg – mass of the clay ball.

According to the task statement the football and the clay ball stopped after the collision, thus due to the law of conservation of momentum we get

p_f=p_c

p_f=m_c v_c

v_c=p_f/m_c =(1.344 (kg⋅m)/s)/(1.2 kg)=1.12 m/s